What is the mass of
0.5 mole of water molecules?
What is the mass of
0.5 mole of water molecules?
 

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Mole of water molecule = 0.5 mole
Molar mass of water molecules = 2 × 1 + 16 = 18 g
Mass of H2O = 18 × 0.5 = 9 g

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Mass of 1 mole water = 18 g
No. Of moles = 0.5
Mass of 0.5 moles = 18 × 0.5 = 9 g
 
Hope it helps!
Thank you ✌️